# Towards a Robust Transport Network With Self-adaptive Network Digital Twin

### v1

@@ 3.1 Physical twin
As the PTwin, we considered a transport network topology modeled as a 2-tuple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. In this graph, $\mathcal{V}$ denotes the set of nodes, each representing a network switch, while $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ defines the set of edges (i.e., links) that connect pairs of nodes. Within this network, a traffic generator can produce a set of flow traffic instances denoted as =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{t},\dots,f_{T}\}$, with a total of $T$ flows. Each flow instance consumes network resources, such as link bandwidth, by applying variable load intensities along a specific path composed of links =$\mathcal{E}^{\prime}=\{e_{1},e_{2},\dots,e_{j},\dots,e_{J}\}$, where $J$ denotes the flow length. In this case, we define each flow instance as a tuple $f_{t}=(\mathbf{x}_{f_{t}},O,D,\mathbf{j})$, where $\mathbf{x}_{f_{t}}\in\mathbb{R}^{5}$ is a feature vector that characterizes the flow’s behavior. In this sense, each feature vector $\mathbf{x}_{f_{t}}$ is associated, by a mapping rule $h$, to a real number $y$ defined as the total per-flow delay, such that $h:\mathbf{x}_{f_{t}}\rightarrow y$. Furthermore, $O$ and $D$ represent the source and destination nodes ($\{O,D\}\in\mathbb{Z}^{+}$); and $\mathbf{j}$ is a vector of link indices that the $t$-th flow traverses. Similarly, each link instance is defined as a tuple $e_{j}=(\mathbf{x}_{e_{j}},\mathbf{i})$, where $\mathbf{x}_{e_{j}}\in\mathbb{R}^{2}$ is a feature vector describing the link’s characteristics, and $\mathbf{i}$ is a vector of flow indices representing the flows traversing that $j$-th link. The complete list of features for both flow and link instances is provided in Table 2. Furthermore, the average traffic rate $x_{t}\in\mathbf{x}_{f_{t}}$ at time $t$ is also described by a probability distribution $P_{t}$, such that $x\sim P_{t}$.
@@ 3.2 Database management and SDN controller
The database used for the decision-making process is divided into three storage types. The first one, called the traffic database, is related to the data stream that was filtered after the traffic generator created it and will serve as input for the VTwin model to perform the necessary delay prediction, i.e., the total per-flow delay label is unavailable. The second type of database is related to the examples that should be used to retrain our VTwin model, and it is called labeled database, as we manipulated the ground truth delay to update the VTwin model weights when necessary. The last storage refers to the weights database, aimed to store the historical VTwin model weights during the NDT operation.
@@
To store these raw data, we are also considering data collection functions to discard unnecessary packet-level traffic data, such as the Internet Protocol (IP) addresses of the server and client, or transmission time timestamps. Therefore, we aim to store only quality of service (QoS) related features that we can extract from the PTwin environment and that are useful for the VTwin prediction task. It is important to note that we assume that the data required for the retraining process has already been collected, labeled, and stored in our database. Furthermore, issues related to synthetic data generation or real-time data collection for retraining are considered outside the scope of this work.
@@
The interfaces employed in the proposed architecture (Figure 1) enable the indirect exchange of control data between both twins, with the SDN controller serving as the network operating system [42]. To allow this interaction, we utilize the SDN controller as the central bridging tool, leveraging its available application programming interface (API). Specifically, another feature stored in the traffic and labeled database consists of the vectors $\mathbf{i}$ and $\mathbf{j}$, which represent, respectively, the links traversed by a given flow and the flows that have traversed a given link. Consequently, one of the critical functions of the interface module is to work with the data collection’s path collector, which identifies and records the routes taken between each source–destination node pair. The path collector operates based on a specific node identifier, which is a 16-character hexadecimal string derived from the PTwin. Therefore, given that flow traffic was initialized between two pairs of nodes, it is possible to consult, by the controller REpresentational State Transfer (REST) API, the path of each flow. In this case, it is assumed that the SDN controller operates in a reactive forwarding way, that is, considering a forwarding decision and route establishment on demand, working only when a packet should be sent. Moreover, the shortest path defines the routing table of each switch in the network.
@@ 3.3 Virtual twin
Considering the adopted NDT architecture, the VTwin assumes a role in mimicking key aspects of its physical counterpart. In this sense, the VTwin is based on a GNN, a type of deep learning model. Accordingly, the VTwin is responsible for abstracting the most complex organization of the transport network and synthesizing its key aspects to evaluate the interest QoS parameter, keeping a certain level of fidelity in the physical-virtual relationship. For the objective of estimating the total flow delay, it is essential to model the relation between the flow traffic applied to propagate information through the wired medium and the links used for it, as shown in the generic graph in Figure 2. This modeling step is fundamental, as the VTwin input is not the original graph representing the transport network but rather the hypergraph derived from it. The relation between the flows and links of Figure 2 creates an essential circular dependency that is explored to generate the heterogeneous graph used as input to the VTwin model. In this case, we can see that flow $f_{1}$ depends on links $e_{1}$, $e_{2}$, and $e_{5}$. Similarly, the flow $f_{2}$ has a dependency on the links $e_{3}$, $e_{4}$, and $e_{5}$. Therefore, these relationships generate the VTwin input graph depicted in Figure 3.
Figure 2: Links and flows relation in a generic network topology.
@@
Regarding the architecture of the VTwin, we have a GNN based on modified RouteNet-Fermi with a self-attention mechanism [28], as illustrated in Figure 4. This model processes flow and link characteristics through message passing across two gated recurrent unit (GRU)-based layers. The aggregated messages are used by a readout function defined by a multilayer perceptron (MLP) to compute occupancy and estimate the total per-flow delay.
Figure 4: Flow and link features from generic topology processed by the VTwin model.
The embedding initialization has the purpose of processing the node features $\mathbf{x}_{f_{t}}$ and $\mathbf{x}_{e_{j}}$ from a heterogeneous graph obtained from the relationship between the links and the flows, that is, the $f_{t}$ flows that passed by several links and the $e_{j}$ links that received each flow [28]. Since this graph has two types of nodes with different feature vectors, specific embedding initialization functions are used to convert the features of each node type from their original dimension to high-dimensional space $\mathbf{h}_{f_{t}}\in\mathbb{R}^{m},\mathbf{h}_{e_{j}}\in\mathbb{R}^{n}$, where $m$ and $n$ are model hyperparameters, respectively, the size of the flow and link embedding vectors.
@@
$\mathbf{h}_{f_{t}}=\text{MLP}(\mathbf{x}_{f_{t}},\theta_{f_{t}}),$ (1)
@@
$\mathbf{h}_{e_{j}}=\text{MLP}(\mathbf{x}_{e_{j}},\theta_{e_{j}}),$ (2)
@@
In the message-passing phase, these high-dimensional vectors are processed iteratively during $k\leq K$ (which $K$ is another model hyperparameter) time steps by the message function $\psi(\cdot)$ and updated by functions $\phi_{\textrm{f}}(\cdot)$, $\phi_{\textrm{l}}(\cdot)$, as defined as follows [43]:
$\mathbf{h}_{f_{t}}^{k+1}=\phi_{\textrm{f}}\left(\mathbf{h}_{f_{t}}^{k},\psi(\mathbf{h}_{f_{t}}^{k},\mathbf{h}_{e_{j}}^{k})\right),$ (3)
@@
$\mathbf{h}_{e_{j}}^{k+1}=\phi_{\textrm{l}}\left(\mathbf{h}_{e_{j}}^{k},\bigoplus_{p\in\mathcal{N}_{e}}a(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1})\psi(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1})\right),$ (4)
where $\bigoplus$ is an aggregation function, $a(\cdot)$ the attention score, and $\mathcal{N}_{e}$ is the neighborhood of $e$-th link node. Moreover, this model treats the message function as a GRU layer and the auxiliary assignment operation as an update function.
The output of the message-passing module computes the occupancy $\mathcal{O}$ through a readout function, implemented using an MLP with two hidden layers. The MLP produces a single-neuron output that yields the occupancy rate for each link in the network, as defined by the following equation:
$\mathcal{O}=\text{MLP}(\mathbf{h}_{f_{t}}^{k+1},\theta_{\textrm{o}}),$ (5)
where $\theta_{\textrm{o}}$ are the learnable parameters of the readout MLP layer, considering that a flow set passes through $J$ links, generating the occupancy mentioned above $\mathcal{O}$, the predicted total per-flow delay $\hat{y}$ is evaluated by:
$\hat{y}=\sum_{j=0}^{J}\left(\dfrac{\mathcal{O}_{j}}{\mathcal{E}_{j}^{\prime}}\right).$ (6)
@@ 3.4 NDT Synchronization
A well-defined LCM provides the NDT platform with the essential capabilities to function effectively in a production environment. It defines how the NDT should behave under various conditions, especially during the operational phase [14], whether those conditions are routine or exceptional. Consequently, all LCM phases can encompass different modules of the architecture presented in Figure 1, and each of them has a specific role and lifespan to ensure the correct operation of the NDT tasks [14]. Since each phase has a different lifespan, some modules shown in Figure 1 may not exist in particular phases. For instance, during the planning stage of the VTwin, the PTwin is not yet required. At this point, the VTwin can be trained using pre-existing datasets [12], considering related features from the target network environment, eliminating the need for real-time data integration. As a result, only the VTwin and selected components of the data management module are active. This phase closely resembles the process of training an AI model for predictive tasks using offline data, where the focus is on optimizing model parameters and hyperparameters based on generalization performance. Noteworthy is that the early stages of NDT for the transport network, as seen in the related works, had a specific focus in this phase.
As the NDT lifecycle progresses, the training phase begins to incorporate physical components of the SDN infrastructure, including the interfaces connected to the SDN controller. In this context, training refers to the process of evaluating the VTwin prior to its deployment in a production environment. This phase enables the identification and mitigation of potential critical issues that could impact the proper functioning of the NDT platform. Finally, in the operational phase, we employ two submodules that monitor the synchronization states of the NDT: concept drift detector and VTwin model management. The relationship in terms of input and output during the NDT operational phase is defined by Figure 5. These submodules are essential as they ensure the NDT robustness against traffic variability after NDT planning and deployment, which is intimately associated with the degree of synchronization between the VTwin and PTwin [15, 33]. This twinning rate will determine the frequency at which the NDT will start the process of synchronizing the network states of the VTwin [37].
Figure 5: Sequence of steps in the operational phase to synchronize the PTwin states in the VTwin.
@@
In a data-driven supervised learning approach, the synchronization process is associated with a retraining operation, which can be triggered based on the input or output of the neural network model. This choice is essential, as the synchronization between the VTwin and PTwin will differ significantly depending on the target variable considered in the concept drift detector process. In the first case, synchronization is driven by the distribution pattern of each input feature, such as the flow of traffic passing through a set of links. In this sense, when decisions regarding concept drift are made solely based on the input data, the approach is referred to as a distribution-based method [44]. In contrast, when the correct labels are also considered, the method is classified as performance-based [44]. In the latter approach, in the context of this problem, synchronization depends on the label, specifically the total per-flow delay. However, using this approach, under realistic conditions, detecting the aging of a neural network model is not straightforward, as the ground truth delay is not readily available for direct comparison with VTwin’s predictions. Therefore, our synchronization approach employs a univariate distribution-based method that monitors the flow traffic rate to detect concept drift during the VTwin operation, thereby eliminating the need to rely on ground truth delay to trigger the synchronization process through VTwin retraining. In operational terms, to reach these synchronization steps, we need to pass through different NDT modules, as defined by the flowchart in Figure 5.
@@
The process starts with user traffic originating from the transport network. This raw traffic data is captured by the relevant network functions and stored in the traffic database. The data collector then extracts a key feature for use by the concept drift detector: the average traffic rate of each flow. This feature is selected because it effectively captures the primary characteristic of user traffic, its average throughput. Thereby, using a univariate covariate-based method as a concept drift detector, we can identify changes in the input distribution of features over time when $P_{t}(x_{f_{t}})\neq P_{t+w}(x_{f_{t+w}})$, where $t$ represents a given time instant, $\{x_{f_{t}}\in\mathbf{x}_{f_{t}},x_{f_{t+w}}\in\mathbf{x}_{f_{t+w}}\}$ is a feature in time instant $t$ and $t+w$, in this case, the average traffic rate, and $w$ are the sizes of the time window. In this context, our concept drift detector serves as a monitoring agent that detects changes in traffic patterns as incoming traffic is processed. This process allows us to indirectly estimate when the total per-flow delay distribution will begin to shift, signaling the expected degradation in VTwin performance and triggering retraining.
@@
When concept drift is detected, the concept drift detector notifies the VTwin model management submodule with a warning to start retraining using data already available in the labeled database. Upon the start of the retraining process, the VTwin model management locks the trigger to prevent the drift detector from issuing duplicate warnings for the same drift event previously addressed until the training is finished. Upon completion of the retraining process, the VTwin model management sends a control message to the VTwin, archives the previously trained model weights in a dedicated database, and updates the VTwin model with the newly learned weights.

### v2

@@ 3.1 Physical twin
As the PTwin, we considered a transport network topology modeled as a 2-tuple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. In this graph, $\mathcal{V}$ denotes the set of nodes, each representing a network switch, while $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ defines the set of edges (i.e., links) that connect pairs of nodes. Within this network, a traffic generator can produce a set of flow traffic instances denoted as $\mathcal{F}=\{f_{1},f_{2},\dots,f_{t},\dots,f_{T}\}$, with a total of $T$ flows. Each flow instance consumes network resources, such as link bandwidth, by applying variable load intensities along a specific path composed of links $\mathcal{E}^{\prime}=\{e_{1},e_{2},\dots,e_{j},\dots,e_{J}\}$, where $J$ denotes the flow length. We also define each flow instance as a tuple $f_{t}=(\mathbf{x}_{f_{t}},O,D,\mathbf{j})$, where $\mathbf{x}_{f_{t}}\in\mathbb{R}^{4}$ is a feature vector that characterizes the flow’s behavior. In this sense, each feature vector $\mathbf{x}_{f_{t}}$ is associated with a mapping rule $h$ to a real number defined as the interested QoS metric, such that $h:\mathbf{x}_{f_{t}}\rightarrow\{y^{(q)}\}$, where $q\in\{\text{delay},\textrm{jitter}\}$, and such that $y^{\textrm{delay}}$ and $y^{\textrm{jitter}}$ represent the per-flow delay and jitter, respectively. Furthermore, $O$ and $D$ represent the source and destination switches ($\{O,D\}\in\mathbb{Z}^{+}$); and $\mathbf{j}$ is a vector of link indices that the $t$-th flow traverses. Similarly, each link instance is defined as a tuple $e_{j}=(\mathbf{x}_{e_{j}},\mathbf{i})$, where $\mathbf{x}_{e_{j}}\in\mathbb{R}^{2}$ is a feature vector describing the link’s characteristics, and $\mathbf{i}$ is a vector of flow indices representing the flows traversing that $j$-th link. The complete list of features for both flow and link instances is provided in Table 3. Among these features, special attention should be given to the propagation delay, which was converted from a link-level feature to a flow-level feature by summing the propagation delays $p_{j}$ of the $j$-th links along the path $\mathcal{E}^{\prime}$. Furthermore, the average traffic rate $x\in\mathbf{x}_{f_{t}}$ at time $t$ is modeled as a random variable following a probability distribution $P_{t}$, such that $x\sim P_{t}$.
@@ 3.2 Database management and SDN controller
The database used for the decision-making process is divided into three types of storage. The first one, called the traffic database, is related to the data stream that was filtered after the traffic generator created it and will serve as input for the VTwin model to perform the necessary QoS prediction; i.e., the total per-flow delay or jitter labels are unavailable. The second type of database is related to the examples that should be used to retrain our VTwin model, and it is called a labeled database, as we manipulated the ground truth delay or jitter to update the VTwin model weights when necessary. The last storage refers to the weights database, which is aimed at storing the historical VTwin model weights during the NDT operation. In summary, the primary roles of each database are depicted in Figure 2.
Figure 2: Primary roles of each database in the NDT data management module.
@@
To store this raw data, we are also considering data collection functions to discard unnecessary packet-level traffic data, such as the Internet Protocol (IP) addresses of the server and client and transmission timestamps. Therefore, we aim to store only QoS related features that we can extract from the PTwin environment and that are useful for the VTwin prediction task. It is important to note that we assume the data required for the retraining process has already been collected, labeled, and stored in our database. Furthermore, issues related to synthetic data generation or real-time data collection for retraining are considered outside the scope of this work.
@@
The interfaces employed in the proposed architecture (Figure 1) enable the indirect exchange of control data between both twins, with the SDN controller serving as the network operating system [Kreutz2015]. This type of interaction is similar to that of a controller-based CPS [molina2018software, salazar2019, kathiravelu2019sd], which also leverages SDN controller interfaces to translate high-level network policies into control instructions executed at the physical layer. With this context, we utilize the SDN controller as the central bridging tool, leveraging its available application programming interface (API). Specifically, another feature stored in the traffic and labeled database consists of the vectors $\mathbf{i}$ and $\mathbf{j}$. Consequently, one of the critical functions of the interface module is to work with the data collection’s path collector, which identifies and records the routes taken between each source–destination node pair. The path collector operates based on a specific node identifier, which is a 16-character hexadecimal string derived from the PTwin. Therefore, given that flow traffic was initialized between two pairs of nodes, it is possible to consult the path of each flow through the controller REpresentational State Transfer (REST) API. In this case, it is assumed that the SDN controller operates in a reactive forwarding manner; that is, it considers forwarding decisions and route establishment on demand, working only when a packet needs to be sent. Proactive forwarding can also be considered in this context by the SDN controller, since this configuration is transparent to VTwin (VTwin only needs to know the path that a given flow traverses). Moreover, for different source destination switch pairs, we assume that the shortest path defines the routing table between them.
@@ 3.3 Virtual twin
Considering the adopted NDT architecture, the VTwin assumes a role in mimicking key aspects of its PTwin. In this sense, the VTwin is based on a GNN, a type of deep learning model. Accordingly, the VTwin is responsible for abstracting the most complex organization of the transport network and synthesizing its key aspects to evaluate the interest QoS parameter while maintaining a certain level of fidelity in the physical-virtual relationship. For the objective of estimating the total per-flow delay and jitter, it is essential to model the relationship between the flow traffic applied to propagate information through the wired medium and the links used for this purpose, as shown in the generic graph in Figure 3. This modeling step is fundamental, as the VTwin input is not the original graph representing the transport network, but rather the hypergraph derived from it. The relationship between the flows and links of Figure 3 creates an essential circular dependency that is explored to generate the heterogeneous graph used as input for the VTwin model. In this case, we can see that flow $f_{1}$ depends on links $e_{1}$, $e_{2}$, and $e_{5}$. Similarly, the flow $f_{2}$ has a dependency on the links $e_{3}$, $e_{4}$, and $e_{5}$. Therefore, these relationships generate the VTwin input graph depicted in Figure 4.
Figure 3: Links and flows relationship in a generic network topology.
@@
Regarding the architecture of the VTwin, we have a GNN based on a modified RouteNet-Fermi with a self-attention mechanism [Modesto2024], as illustrated in Figure 5. This model processes flow and link characteristics through message passing across two gated recurrent unit (GRU)-based layers. The aggregated messages are used by a readout function defined by a multilayer perceptron (MLP) to compute the parameters used to estimate QoS metrics.
Figure 5: The adopted GNN architecture takes as input both flow and link features from the heterogeneous graph. It is composed of three modules: embedding initialization, message passing, and readout. Its output consists of QoS metrics such as delay and jitter.
The embedding initialization aims to process the node features $\mathbf{x}_{f_{t}}$ and $\mathbf{x}_{e_{j}}$ from a heterogeneous graph obtained from the relationship between the links and the flows; that is, the $f_{t}$ flows that passed through several links and the $e_{j}$ links that received each flow [Modesto2024]. Since this graph has two types of nodes with different feature vectors, specific embedding initialization functions are used to convert the features of each node type from their original dimensions to high-dimensional space $\mathbf{h}_{f_{t}}\in\mathbb{R}^{m},\mathbf{h}_{e_{j}}\in\mathbb{R}^{n}$, where $m$ and $n$ are model hyperparameters, respectively, representing the sizes of the flow and link embedding vectors.
@@
$\mathbf{h}_{f_{t}}=\text{MLP}\left(\mathbf{x}_{f_{t}},\theta_{f_{t}}\right),$ (1)
@@
$\mathbf{h}_{e_{j}}=\text{MLP}\left(\mathbf{x}_{e_{j}},\theta_{e_{j}}\right),$ (2)
@@
In the message-passing phase, these high-dimensional vectors are processed iteratively during $k\leq K$ (which $K$ is another model hyperparameter) time steps by the message function $\psi(\cdot)$ and updated by functions $\phi_{\textrm{f}}(\cdot)$, $\phi_{\textrm{l}}(\cdot)$, as defined as follows [Bronstein2021GeometricDL]:
$\mathbf{h}_{f_{t}}^{k+1}=\phi_{\textrm{f}}\left(\mathbf{h}_{f_{t}}^{k},\psi\left(\mathbf{h}_{f_{t}}^{k},\mathbf{h}_{e_{j}}^{k}\right)\right),$ (3)
@@
$\mathbf{h}_{e_{j}}^{k+1}=\phi_{\textrm{l}}\left(\mathbf{h}_{e_{j}}^{k},\bigoplus_{p\in\mathcal{N}_{e}}a\left(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1}\right)\psi\left(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1}\right)\right),$ (4)
where $\bigoplus$ is an aggregation function, $a(\cdot)$ is the attention score, and $\mathcal{N}_{e}$ is the neighborhood of the $e$-th link node. Moreover, this model treats the message function as a GRU layer, and the auxiliary assignment operation as an update function. The output of the message-passing is used by the readout function to calculate the parameter $\mathcal{O}$, implemented using an MLP with two hidden layers. This MLP produces a single-neuron output, as defined by the following equation:
$\mathcal{O}=\text{MLP}\left(\mathbf{h}_{f_{t}}^{k+1},\theta_{\textrm{o}}\right),$ (5)
where $\theta_{\textrm{o}}$ are the learnable parameters of the readout MLP layer.
Considering that a flow set passes through $J$ links, in the delay prediction task, this parameter $\mathcal{O}$ can be interpreted as queue occupancy, such that the predicted total per-flow delay $\hat{y}^{(\text{delay})}$ is evaluated by:
$\hat{y}^{(\textrm{delay})}=\underbrace{\sum_{j=0}^{J}\left(\dfrac{\mathcal{O}_{j}}{\mathcal{E}_{j}^{\prime}}\right)}_{\text{Queue delay}}\hskip 4.26773pt+\underbrace{\sum_{j=0}^{J}p_{j}.}_{\text{Propagation delay}}$ (6)
Although no explicit variations in queue configuration (such as buffer size) are introduced, the predicted total per-flow delay $\hat{y}^{\text{delay}}$ in Eq. (6) inherently includes the effect of queuing delay, derived using the output from Eq. (5), as the queue states implicitly affect each flow’s state [ferriol2023routenet].
Finally, the predicted total per-flow jitter is directly evaluated using the output of readout function, as defined by:
$\hat{y}^{(\textrm{jitter})}=\sum_{j=0}^{J}\left(\dfrac{\mathcal{O}_{j}}{\mathcal{E}_{j}^{\prime}}\right).$ (7)
@@ 3.4 NDT Synchronization
A well-defined LCM provides the NDT platform with the essential capabilities to function effectively in a production environment. It defines how the NDT should behave under various conditions, especially during the operational phase [Kherbache2021], whether those conditions are routine or exceptional. Consequently, all LCM phases can encompass different modules of the architecture presented in Figure 1, each of which has a specific role and lifespan to ensure the correct operation of the NDT tasks [Kherbache2021]. Since each phase has a different lifespan, some modules shown in Figure 1 may not exist in particular phases, as depicted in Figure 6.
Figure 6: The NDT LCM comprises three phases: planning, training, and operational. Initially, no module is activated. During the planning phase, only the data management and VTwin modules are enabled (purple blocks). In the training phase, interaction with the PTwin by the SDN interfaces and the use of a network application are required (light magenta blocks). Finally, in the operational phase, all modules are activated, and NDT synchronization (light blue block) operates to maintain adequately the performance of the VTwin.
For instance, during the planning stage of the VTwin, the PTwin is not yet required. At this point, the VTwin can be trained using pre-existing datasets [bariah2024interplay], considering related features from the target network environment and eliminating the need for real-time data integration. As a result, only the VTwin and selected components of the data management module are active. This phase closely resembles the process of training an AI model for predictive tasks using offline data, where the focus is on optimizing model parameters and hyperparameters based on generalization performance. Noteworthy is that the early stages of NDT for the transport network, as seen in the related works, had a specific focus during this phase. As the NDT lifecycle progresses, the training phase begins to incorporate the physical components of the SDN infrastructure, including the interfaces connected to the SDN controller. In this context, training refers to the process of evaluating the VTwin prior to its deployment in a production environment. This phase enables the identification and mitigation of potential critical issues that could impact the proper functioning of the NDT platform. Finally, in the operational phase, we employ two submodules that monitor the synchronization states of the NDT: the concept drift detector and VTwin model management. These submodules are essential as they ensure the NDT robustness against traffic variability after NDT planning and deployment, which is intimately associated with the degree of synchronization between the VTwin and PTwin [Alghamdi2024, liu2024digital]. This twinning rate will determine the frequency at which the NDT will start the process of synchronizing the network states of the VTwin [jones2020].
@@
In a data-driven supervised learning approach, the synchronization process is associated with a retraining operation [Bagui2025] that can be triggered based on the input or output of the neural network model, as depicted in Figure 7(a) and 7(b), respectively. In this context, concept drift can be detected using input features such as the average traffic rate or output features such as per-flow delay and jitter. These features may be analyzed either individually or jointly by the concept drift detector. When considered jointly, the detection relies on a multivariate concept drift detector [lukats2025benchmark], whereas analyzing each feature independently corresponds to a univariate approach.
(a) Concept drift based on the changes of average traffic rate
(b) Concept drift based on the changes of per-flow delay or jitter
Figure 7: Types of data sources that can trigger a retraining process in a data-driven NDT.
@@
This choice is essential, as the synchronization between the VTwin and PTwin will differ significantly depending on the target variable considered in the concept drift detection process. In the first case, synchronization is driven by the distribution pattern of each input feature, such as the network traffic of a flow passing through a set of links. In this sense, when decisions regarding concept drift are made solely based on the input data, the approach is referred to as a distribution-based method [Bayram2022]. In contrast, when the correct labels are also considered, the method is classified as performance-based [Bayram2022]. In the latter approach, within the context of this problem, synchronization depends on the label, specifically on the total per-flow delay or jitter. However, using this approach under realistic conditions, detecting the aging of a neural network model is not straightforward, as the ground truth of the QoS metric is not readily available for direct comparison with VTwin’s predictions. Therefore, our synchronization approach employs a univariate distribution-based method that monitors the average flow traffic rate to detect concept drift during the VTwin operation, thereby eliminating the need to rely on ground truth delay or jitter to trigger the synchronization process through VTwin retraining. In operational terms, to reach these synchronization steps, we need to pass through different NDT modules, as defined by the flowchart in Figure 8.
Figure 8: Sequence of steps in the operational phase to synchronize the VTwin states based on the current PTwin.
@@
The process starts with user traffic originating from the transport network. This raw traffic data is captured by the relevant network functions and stored in the traffic database. The data collector then extracts a key feature for use by the concept drift detector: the average traffic rate of each flow. This feature is selected because it effectively captures the primary characteristic of user traffic: its average throughput. Thereby, using a univariate covariate-based method as a concept drift detector, we can identify changes in the input distribution of features over time when $P_{t}(x_{f_{t}})\neq P_{t+w}(x_{f_{t+w}})$, where $t$ represents a given time instant, $\{x_{f_{t}}\in\mathbf{x}_{f_{t}},x_{f_{t+w}}\in\mathbf{x}_{f_{t+w}}\}$ is a feature in time instant $t$ and $t+w$, in this case, the average traffic rate, and $w$ are the sizes of the time window. In this context, our concept drift detector serves as a monitoring agent that detects changes in traffic patterns as incoming traffic is processed. This process allows us to indirectly estimate when the total QoS metric distribution will begin to shift, signaling the expected degradation in VTwin performance and triggering retraining.
@@
When a concept drift is detected, the concept drift detector notifies the VTwin model management submodule with a warning to start retraining using data that is already available in the labeled database. Upon the start of the retraining process, the VTwin model management locks the trigger to prevent the drift detector from issuing duplicate warnings for the same drift event that has been previously addressed until the training is finished. Upon completion of the retraining process, the VTwin model management sends a control message to the VTwin, archives the previously trained model weights in a dedicated database, and updates the VTwin model with the newly learned weights.